# ANVIL: Anomaly-based Vulnerability Identification without Labelled Training Data

### v1

@@ III Evaluation
  • RQ1: How do different anomaly scoring methods affect the ANVIL’s vulnerability detection capability?
@@ III-A Experiment Setup
All experiments were conducted using LLMs and tokenizers from the Huggingface library [18], on a machine equipped with an Intel Xeon 4509Y CPU and an Nvidia H100 GPU with 80GB of HBM. For all LLMs, we used a temperature value of zero during code generation to prevent randomness.
We use the Magma [19] fuzzing benchmark to evaluate ANVIL’s vulnerability detection ability. The Magma benchmarks contain 138 real-world CVEs from 9 widely-used open source projects in C/C++ (libpng, libtiff, libsndfile, libxml2, poppler, openssl, sqlite3, php, lua), containing approximately 240,000 lines and 5,227 functions in total. Additionally, Magma includes both CVE (vulnerable) and patched (non-vulnerable) versions of these projects. In the vulnerable version, there are 256 lines labelled as vulnerable, of which 251 lines are located in 158 functions, while the remaining 5 lines are not associated with any function. One reason we use Magma is that as mentioned by Chen et al. [12], other popular vulnerability datasets such as BigVul [20] and CVEFixes [21] suffer from high mislabeling rates. These datasets usually consider all deleted code in a bug-fixing commit as vulnerable, potentially including irrelevant content and making the evaluation inaccurate. In contrast, the labelling of vulnerable lines in Magma is significantly more reliable because Magma developers provide a Proof of Vulnerability (POV) for every bug, which represents an input that triggers the vulnerability. These POVs are extracted through manual review of the bug reports and source files, as well as utilizing multiple fuzzing campaigns, ensuring all vulnerabilities in Magma are both reachable and exploitable.
@@
Additionally, to address concerns of LLM data leakage, in RQ5, we test our methodology with vulnerabilities unseen by the LLMs. Because each LLM we tested is trained at a different time, it is difficult to find a large set of vulnerabilities that is guaranteed to be unseen by all models. Therefore, we choose to test only the CodeLlama-13B model for leakage experiments because it generally provides the best performance (as shown by having the lowest P-values in Table III). We leave the construction of leakage-free datasets and evaluation of other LLMs for future work. In total, we collected 69 CVEs from 35 different repositories using a script from the CVEFixes project [21]. All collected CVEs were published after December 31, 2023, which is later than knowledge-cutoff of the CodeLlama-13B model we used, ensuring that the CVEs are unseen by the model. Instead of only relying on lines changed by defect-fixing commits, which may contain irrelevant information, we manually inspect each commit to extract only the vulnerable lines of codes. This resulted in a dataset of 60 vulnerable source files for the leakage experiment, containing 156 vulnerable LOC with 144 located within 104 functions and 12 not associated with any specific function.
@@ III-B RQ1: Anomaly Score
To evaluate which scoring function described in Section II-C can best distinguish vulnerable lines as anomalies, we conducted experiments using the CodeLlama-13B model on the Magma dataset. Specifically, we examine the distributions generated by each anomaly score for each line type (i.e. vulnerable and non-vulnerable) and evaluate the ability of each anomaly score to discriminate between the vulnerable and non-vulnerable distributions. We utilize the Mann-Whitney U test to calculate a P-value between the two distributions for each score, where a lower P-value indicates a higher statistical significance that the LLM generates different score distributions between the two line types. We have adopted the commonly utilized threshold of $\alpha=0.05$ to determine if the two line types can be significantly distinguished, providing a confidence level of at least 95%.
We employed the FIM task, using 256 vulnerable LOC from the Magma dataset as vulnerable samples. Considering the impracticality of using over 250,000 non-vulnerable lines due to vast dataset size and time constraints, we balanced the dataset by randomly sampling 256 non-vulnerable lines for a comparative analysis. Additionally, we evaluated the vulnerability-anomaly detection capability across various fixed context sizes ranging from [500, 400, 300, 200, 100], split evenly between the prefix and the suffix surrounding the masked line. This range was selected based on observed performance metrics: context sizes smaller than 100 led to significant performance drops due to insufficient information for FIM (as illustrated in Figure 3), while sizes larger than 500 often caused out-of-memory errors on the GPU, potentially compromising evaluation accuracy.
Figure 3: Median anomaly score on various functions
For each reconstructed line of code, we calculated the anomaly scores using the aforementioned three scoring functions. Figure 3 displays three graphs, one per scoring function, with each graph showing the median anomaly scores across various fixed context sizes. We chose to plot the median value instead of the mean because, during our experiments, we observed several outliers in the anomaly scores, which could potentially skew the mean and provide a less accurate representation of the data. We observe a noticeable performance drop when the context size is below 100 lines, but once the performance stabilizes, the results consistently show a discernible gap between vulnerable and non-vulnerable lines, reinforcing the notion that the LLM treats these two categories differently. Among the scoring functions, the gap is larger when using the hybrid function compared to only the loss function.
We then computed P-values between the distributions of scores for vulnerable and non-vulnerable lines for each scoring function and context size pair, as shown in Table I. For each context size column, we signal in bold, the scoring function that produced the lowest P-value. We can see that the combined scoring function (Equation 1) most frequently (i.e. largest number of wins) results in the smallest P-values between the two line types.
TABLE I: P-values using different scoring functions
@@
Function 100 200 300 400 500 Wins
$\delta_{1}$ 7.1e-4 2.2e-4 2.7e-5 1.4e-4 0.0025 1
$\delta_{2}$ 4.7e-5 0.0103 6.8e-5 0.0011 0.0059 1
$\delta$ 3.3e-4 4.9e-4 9.4e-6 9.0e-5 0.0020 3
@@ III-C RQ2: Model Architecture and Model Size
While RQ1 uses CodeLlama-13B as a representative model, in this section, we investigate whether this capability of distinguishing vulnerabilities as anomalous generalizes to other LLMs. To this end, we conduct the FIM workload on various open-source code LLMs. Open-source models, unlike closed-source ones such as OpenAI’s GPT-3.5/4 series [22], provide not only the final generated tokens but also allow access to intermediate results like loss values and GPU usage, offering greater transparency and adaptability for research needs. We select the top-ranked base models from Huggingface’s Big Code Models Leaderboard as of May 1st, 2024 [23], focusing on those that support FIM code generation. To manage GPU memory efficiently, we restricted model parameters to fewer than 15 billion. Furthermore, during our evaluation, we observed that the StarCoder2-15B model exhibited poor performance on the FIM task due to an implementation bug highlighted in its original paper [24]. We thus substituted it with the StarCoderBase-15B model [25].
@@
Model Name Model Size Training data size  [^1] Approximated using 4 characters per token and 70 characters per LOC. Publication Time
@@
Therefore, the evaluation involves four models, which are detailed in Table LABEL:Tab:model_list [26, 27, 25]. Our experiments applied the FIM task across these LLMs, assessing their median anomaly score under various fixed context sizes, as depicted in Figure 4. These figures reveal common patterns across the LLMs: 1) All models showed a drop in performance with inadequate context sizes, typically under 100 lines; 2) After LLMs acquire enough context information, their anomaly scores become stable, except for slight fluctuations observed in the CodeQwen-7B model with non-vulnerable lines; and 3) Significant gaps were observed between the scores for vulnerable and non-vulnerable lines across all models.
To precisely determine whether these LLMs can distinguish between the two line types, we calculated P-values for contexts larger than 100 lines, as shown in Table III. We can see that, all P-values consistently fell below the $\alpha=0.05$ threshold across all models and sizes, highlighting the effectiveness of all LLMs in distinguishing vulnerable code as anomalous.
TABLE III: P-values using different LLMs
@@
Function 100 200 300 400 500
CodeLlama-13B 3.3e-4 4.9e-4 9.4e-6 9.0e-5 0.0020
CodeLlama-7B 0.0012 0.0019 1.2e-4 0.0026 0.0076
CodeQwen 3.9e-4 6.0e-4 3.6e-4 0.0026 0.0078
StarCoderBase 0.0010 0.0034 2.9e-4 1.4e-4 4.2e-4
When comparing different sizes of the same model, particularly CodeLlama-13B and CodeLlama-7B, the larger model produced lower P-values for all context sizes, indicating that it has greater discriminating ability. However, this improvement is not without trade-offs. The larger CodeLlama model requires significantly more GPU memory for inference. On our H100 GPU, we discovered three samples which triggered CUDA out-of-memory errors when using large fixed context sizes; specifically, for 400 and 500 line contexts, one sample and two samples respectively trigger this error due to long line lengths. In contrast The 7B parameter model had a peak memory usage of 52GB with the same workload. This reduction in memory usage makes the smaller model a better choice when GPU memory is limited. While we have only conducted experiments on the CodeLlama models, these results are promising, and suggest that vulnerability detection accuracy should improve with neural model scaling.
@@ III-D RQ3: Adaptive Context Size
While fixed context sizes have demonstrated fairly good performance in anomaly detection, we believe that this approach can be improved for general anomaly and vulnerability discovery. A fixed context length, such as a 150-line prefix and suffix, may not sufficiently encapsulate all relevant code features for vulnerability detection. In cases where the relevant context extends beyond this 300-line window, such as in functions longer than 300 lines, critical contextual information will not be conveyed to the LLM. Conversely, in smaller functions that span only 50 lines, a 300-line context window could introduce irrelevant information. Although our previous findings suggest that this surplus information does not significantly degrade performance, it does lead to unnecessary GPU memory and computational resource usage.
As a result, we propose that utilizing the maximum compound statement (MCS) as context will improve on fixed context sizes. Using the MCS, which is typically the function body, ensures that the model receives semantically relevant context to effectively learn the code’s features without being burdened by irrelevant information that the model might inherently ignore.
After adopting MCS as our context configuration, we conducted the FIM experiment using the CodeLlama-13B model. Comparing against fixed context sizes, as shown in Figure 5, MCS significantly lowered the P-value between vulnerable and non-vulnerable lines to 1.3e-6, a significant improvement compared to the lowest P-value obtained with a fixed 300-line context, which was 9.4e-6. Furthermore, the MCS approach resolved the CUDA out-of-memory issues previously encountered with larger fixed contexts, reducing peak memory usage to 67GB and enhancing energy efficiency. These results illustrate an important insight—while varying fixed context size on a set of code lines has some effect on average discrimination ability, there can be specific lines where an inappropriate context size significantly confuses the model, and a dynamic context size based on code structure is better for discriminating between vulnerable and non-vulnerable code.
Figure 5: P-values using fixed context sizes vs. MCS
@@ III-E RQ4: Line-level Vulnerability Detection
In this section, we compare ANVIL, which does not require any labeled vulnerability data, against two state-of-the-art supervised-learning-based vulnerability detectors: LineVul and LineVD. LineVul [8] is a software vulnerability detector that offers both function-level and line-level detection capabilities. It fine-tunes CodeBert [1] using the BigVul [20] vulnerability dataset. The detection process begins at the function level, where LineVul assesses whether a function contains any vulnerabilities. If a function is deemed vulnerable, LineVul then evaluates each line within that function. It does this by aggregating the attention scores from CodeBert’s transformer layers for each token in a line, resulting in a composite attention score for that line. A higher score means the line is more likely to be vulnerable. The intuition is that, for a vulnerability prediction task, lines with higher attention scores are deemed more critical to the model output, suggesting a greater likelihood of being linked to the root cause of the vulnerability.
@@
LineVD [9], like LineVul, also supports both function-level and line-level detection. LineVD also uses CodeBert to generate code embeddings for each function and its constituent lines. However, it extends its analysis framework by utilizing Joern [28] to create a graph representation for each function. This graph, along with the code embeddings, are processed using a graph attention network (GAT) and a multi-layer perceptron (MLP) to simultaneously detect vulnerabilities at both the function and line levels. This system is trained on the BigVul dataset, similar to LineVul. A distinctive feature of LineVD is its consistency mechanism: it employs element-wise multiplication within its model to ensure that if a function is classified as non-vulnerable (label 0), then all lines within that function are also automatically classified as non-vulnerable.
@@
In this section, all experiments are conducted using the CodeLlama-13B model with the hybrid anomaly scoring function (Equation 1) and MCS contexts. To ensure fairness and comparability in our evaluation, we will use the same metrics as those employed by our baseline models. We evaluate ANVIL vs LineVul and LineVD using two metrics: 1) We approach vulnerability detection as a binary classification task and measure the line-level classification performance using the Area Under the Receiver Operating Characteristic Curve (ROC-AUC). This metric treats each line of code as an independent data sample, assessing the overall classification accuracy without setting a specific threshold. 2) To gauge how effective ANVIL’s anomaly score is as a vulnerability prioritization tool, we evaluate all detectors with a ranking task in which each line within a known vulnerable function is assigned a detector specific score. We then use a Top-5 metric (a standard metric for measuring vulnerability prioritization ability) to compare the detector predictions. This measure assesses whether any of the vulnerable lines are included in the detector’s top-5 most confident predictions. For this metric, each function with known vulnerabilities is treated as a separate data sample.
To enhance the reliability of our results without compromising generalizability, we utilize two mechanisms to filter out lines that are unlikely to contain vulnerabilities: M1) When a masked line is at the very beginning of the context, LLMs often misinterpret this as the start of a source file and inappropriately generate an “#include” statement. To prevent this misclassification from skewing our results, such lines are automatically considered non-vulnerable; and M2) when both the masked and generated outputs in the FIM task are comments, they are automatically classified as non-vulnerable. This exclusion is only applied when both are comments as cases where the model predicts that code should have been present, but a comment existed instead, may be indicative of instances where critical code is missing.

### v2

@@ III Evaluation
  • RQ1: How do different anomaly scoring methods affect ANVIL’s vulnerability detection capability?
@@ III-A Experiment Setup
All experiments were conducted using LLMs and tokenizers from the Huggingface library [wolf_transformers_2020], on a machine equipped with an Intel Xeon 4509Y CPU and an Nvidia H100 GPU with 80GB of HBM. For all LLMs, we used a temperature value of zero during code generation to prevent randomness.
Previous studies [chen_diversevul_2023, ding_vulnerability_2024] have revealed that popular vulnerability datasets, such as Devign [zhou_devign_2019] and CVEFixes [bhandari_cvefixes_2021], often suffer from high mislabelling rates. These datasets may mistakenly label commits unrelated to vulnerabilities as vulnerable, and may also include modifications which are irrelevant to vulnerability patching. To mitigate the influence of such mislabelling issues, we carefully selected real-world vulnerability datasets with high-accuracy labels.
Specifically, we employed three datasets to evaluate ANVIL across different dimensions. First, RQ1 through RQ3 serve as internal evaluations, exploring how ANVIL performs under varying configurations, such as different LLMs and context sizes. For these experiments, we selected the Magma [hazimeh_magma_2020] fuzzing benchmark due to its reliable vulnerability labelling. Specifically, Magma’s labels are backed by a Proof of Vulnerability (PoV) for each bug, representing an input that triggers the vulnerability. These PoVs are extracted through manual reviews of bug reports and source files and complemented by extensive fuzzing campaigns, ensuring that all vulnerable and patched lines are directly relevant to the vulnerabilities. Magma contains 138 real-world CVEs from nine widely-used open-source C/C++ projects (libpng, libtiff, libsndfile, libxml2, poppler, openssl, sqlite3, php, and lua), comprising approximately 240,000 LoC and 5,227 functions. The dataset features both CVE (vulnerable) and patched (non-vulnerable) versions of these projects, with 256 lines labelled as vulnerable.
After identifying the most effective configurations of ANVIL, we conduct a detailed comparison between ANVIL and various supervised-learning-based vulnerability detectors to address RQ4. To address potential overfitting from the previous configuration tuning process (RQ1–RQ3) and to demonstrate ANVIL’s ability to generalize to a broader range of real-world projects, we selected the PrimeVul [ding_vulnerability_2024] dataset, which comprises over 700 open-source C/C++ projects. Despite being significantly larger than Magma, PrimeVul still maintains a low labelling error rate. It achieves more than 90% labelling accuracy through automated validation, which analyzes corresponding vulnerability descriptions in the NVD database. For RQ4, we use the PrimeVul validation and test datasets, which consist of 1,142 vulnerable functions and approximately 47,000 benign functions.
@@
Additionally, to address concerns of LLM data leakage, in RQ5, we test our methodology with vulnerabilities unseen by the LLMs during training. Because each LLM we tested is trained at a different time, it is difficult to find a large set of vulnerabilities that is guaranteed to be unseen by all models. Therefore, we choose to test only the CodeLlama-13B model for leakage experiments because it generally provides the best performance, as per its ROC-AUC in Table III. We leave the construction of leakage-free datasets and evaluation of other LLMs for future work. In total, we collected 100 CVEs from 53 different repositories using a script from the CVEFixes project [bhandari_cvefixes_2021]. All collected CVEs were published between December 31, 2023 to April 20, 2024, which is later than the knowledge-cutoff of the CodeLlama-13B model we used, ensuring that the CVEs are unseen by the model. Instead of only relying on lines changed by defect-fixing commits, which may contain irrelevant information, we manually inspect each commit to extract only the vulnerable lines of code. This resulted in a dataset of 147 vulnerable functions, containing 235 vulnerable LoC and 10,956 non-vulnerable LoC. We also included 3,589 benign functions consisting of 101,173 LoC to reflect the mixed and imbalanced nature of vulnerable and benign code in real-world projects. This dataset is publicly available in our repository to support future research.
@@ III-B RQ1: Anomaly Score
To evaluate which scoring function described in Section II-C can best distinguish vulnerable lines as anomalies, we conducted experiments using CodeLlama-13B on the Magma dataset. Specifically, we examine the distributions generated by each anomaly score for each line type (i.e., vulnerable and non-vulnerable) and evaluate the ability of each anomaly score to discriminate between the vulnerable and non-vulnerable distributions. This evaluation was framed as a binary classification task, with the Area Under the Receiver Operating Characteristic Curve (ROC-AUC) score used as a metric to measure the significance of the distinction between vulnerable and non-vulnerable lines. An ROC-AUC score of 0.5 indicates random guessing, while higher scores reflect a greater discrimination capability.
We employed the FIM task, using 256 vulnerable LOC from the Magma dataset as vulnerable samples. Considering the impracticality of using over 240k non-vulnerable lines due to vast dataset size and time constraints, we balanced the dataset by randomly sampling 256 non-vulnerable lines for a comparative analysis. Additionally, we evaluated the vulnerability-anomaly detection capability across various fixed context sizes ranging from [500, 400, 300, 200, 100, 80, 60, 40, 20, 10, 2] lines, split evenly between the prefix and the suffix surrounding the masked line. We limited the context size to a maximum of 500 lines in our evaluation, as sizes larger than this often caused out-of-memory errors on the GPU during our experiments, potentially compromising evaluation accuracy.
Figure 3: Anomaly scores on various functions
TABLE I: ROC-AUC using different scoring functions
@@
Func 2 10 20 40 60 80 100 200 300 400 500 Wins
$\delta_{EM}$ 52.2% 55.3% 56.1% 58.2% 58.5% 60.1% 59.0% 55.5% 59.2% 57.6% 56.6% 1
$\delta_{loss}$ 53.6% 56.3% 56.0% 59.1% 60.1% 61.4% 60.4% 61.1% 62.5% 61.6% 59.6% 3
$\delta_{hybrid}$ 54.1% 55.4% 56.1% 59.4% 60.5% 61.7% 60.6% 60.4% 62.9% 61.7% 59.5% 8
For each reconstructed line of code, we calculated anomaly scores using the three scoring functions described in Section II-C ($\delta_{loss}$, $\delta_{EM}$, and $\delta_{Hybrid}$). Figure 3 presents three graphs, each corresponding to one scoring function. For the Exact Match (EM) scores, we reported the averaged values due to their binary nature. In contrast, for the Loss and Hybrid scores, we plotted the median values instead of the mean. This decision was informed by our observation of several outliers in these two scoring functions, particularly in non-vulnerable lines where the reconstructed line differed significantly from the masked line, which could skew the mean and provide a less accurate representation of the data. To provide a comprehensive view of the Loss and Hybrid scores, we also included box plots [noauthor_box_2025], with the mean values represented by small triangles.
Across all scoring functions, the median scores reveal a discernible gap between vulnerable and non-vulnerable lines, and this gap stabilizes typically when the context size exceeds 100 lines. This observation reinforces the notion that the LLM processes these two categories differently. Conversely, as the context size decreases below 100 lines, the anomaly scores for both vulnerable and non-vulnerable lines gradually increase, and the gap between the two types shrinks, indicating that the LLM becomes less accurate in reconstructing the masked lines regardless of their categories. This performance decline occurs because smaller context sizes do not provide the model with sufficient information for accurate generation. Among the scoring functions, the Hybrid function exhibited a larger median gap between vulnerable and non-vulnerable lines compared to the Loss function alone. Furthermore, the mean values of the two line types for the Loss score showed a smaller gap compared to those under the Hybrid score, which, despite also being affected by outliers, still displayed a more distinguishable separation between the two line types.
To quantify these findings, we computed ROC-AUC scores for the distributions of vulnerable and non-vulnerable lines for each scoring function and context size pair, as shown in Table I. For each context size, we highlighted in bold the scoring function that achieved the highest ROC-AUC score. The results indicate that the combined scoring function ($\delta_{Hybrid}$, Equation 1) most frequently yields the highest ROC-AUC, achieving the largest number of “wins” across context sizes. In contrast, the Exact Match Accuracy consistently produced the lowest scores for most context sizes, earning only one tied win at a context size of 20 lines.
@@ III-C RQ2: Model Architecture and Model Size
While RQ1 uses CodeLlama-13B as a representative model, in this section, we investigate whether this capability of distinguishing vulnerabilities as anomalous generalizes to other LLMs. To this end, we conduct the FIM workload on various open-source code LLMs. Open-source models, unlike closed-source ones such as OpenAI’s GPT-3.5/4 series [noauthor_chatgpt_nodate], provide not only the final generated tokens but also allow access to intermediate results like loss values and GPU usage, offering greater transparency and adaptability for research needs. We select the top-ranked base models from Huggingface’s Big Code Models Leaderboard as of May 1st, 2024 [noauthor_big_nodate], focusing on those that support FIM code generation. To manage GPU memory efficiently, we restricted model parameters to fewer than 15 billion. Furthermore, during our evaluation, we observed that the StarCoder2-15B model exhibited poor performance on the FIM task due to an implementation bug highlighted in its original paper [lozhkov_starcoder_2024]. We thus substituted it with the StarCoderBase-15B model [li_starcoder_2023].
@@
Model Name Size Training data  [^1] Approximated using 4 characters per token and 70 characters per LOC. Publication
@@
TABLE III: ROC-AUC on different LLMs
@@
LLM 2 10 20 40 60 80 100 200 300 400 500
CodeLlama-13B 54.1% 55.4% 56.1% 59.4% 60.5% 61.7% 60.6% 60.4% 62.9% 61.7% 59.5%
CodeLlama-7B 54.0% 56.4% 56.7% 57.8% 59.7% 60.8% 60.0% 59.8% 61.7% 59.6% 58.7%
CodeQwen 54.4% 55.7% 54.8% 58.8% 60.0% 61.9% 60.3% 60.3% 60.6% 59.1% 58.1%
StarCoderBase 55.8% 55.5% 55.3% 58.1% 58.5% 59.8% 60.0% 59.3% 61.0% 61.7% 61.0%
Therefore, the evaluation involves four models, which are detailed in Table LABEL:Tab:model_list [roziere_code_2024, team_code_2024, li_starcoder_2023]. Our experiments applied the FIM task across these LLMs, assessing their median anomaly score under various fixed context sizes, as depicted in Figure 4. This figure reveals common patterns across the LLMs: 1) All models showed a performance drop in mask reconstruction accuracy with inadequate context sizes; 2) After LLMs acquire enough context information, their anomaly scores become constant; and 3) Significant gaps were observed between the scores for vulnerable and non-vulnerable lines across all models.
To precisely determine whether these LLMs can distinguish between vulnerable and non-vulnerable lines, we calculated ROC-AUC scores for contexts between 2 to 500 lines, as shown in Table III. All scores consistently exceeded the 50% threshold of random guessing, and surpassed 60% once the models had sufficient context window sizes, demonstrating the effectiveness of all models in identifying vulnerable lines as anomalous. Notably, the best performance was achieved by CodeLlama-13B with a fixed context size of 300, reaching an ROC-AUC score of 62.9%.
When comparing different sizes of the same model, particularly CodeLlama-13B and CodeLlama-7B, the larger model achieved slightly higher ROC-AUC scores, particularly when performance stabilized at larger context sizes, indicating that it has greater discriminating ability. However, this improvement is not without trade-offs. The larger CodeLlama model requires significantly more GPU memory for inference. On our H100 GPU, we discovered three samples which triggered CUDA out-of-memory errors when using large fixed context sizes; specifically, for 400 and 500 line contexts, one sample and two samples respectively trigger this error due to long line lengths. In contrast, the 7B parameter model had a peak memory usage of 52GB, out of the 80GB available on the H100 GPU with the same workload. This reduction in memory usage makes the smaller model a better choice when GPU memory is limited. While we have only conducted experiments on the CodeLlama models, these results are promising, and suggest that vulnerability detection accuracy should improve with neural model scaling.
@@ III-D RQ3: Adaptive Context Size
While fixed context sizes have demonstrated fairly good performance in anomaly detection, we propose that utilizing the Adaptive Context (AC) will further improve general anomaly and vulnerability discovery. To evaluate the effectiveness of AC, we conducted AC FIM experiments using the CodeLlama-13B model and compared the results to those obtained with fixed context sizes. As shown in Figure 5, the blue bars represent results using fixed context sizes, with values directly taken from the first row of Table III, while the orange bar represents the result achieved with AC. Adopting AC significantly improves the ROC-AUC score between vulnerable and non-vulnerable lines to 63.7%, surpassing all ROC-AUC scores achieved with fixed context sizes. Additionally, the AC approach resolved the CUDA out-of-memory issues encountered with larger fixed contexts, reducing peak memory usage to 67GB and improving energy efficiency. These results highlight a key insight: while varying fixed context sizes can affect the average discrimination ability, certain lines may significantly confuse the model when an inappropriate context size is used. Allowing the context to not have to be centered around the line being analyzed also reduces the presence irrelevant lines that may confuse the model.color=purple!30color=purple!30todo: color=purple!30david A dynamic context size based on code structure, such as AC, proves more effective in distinguishing between vulnerable and non-vulnerable code.
We further measured the time required to conduct experiments under different context sizes, as shown in Figure 5. The graph reveals a clear trend: with fixed context sizes, the elapsed time increases as the context window enlarges. This outcome is expected, as larger context sizes demand more computational resources, leading to increased contention for hardware resources and, ultimately, slower inference times for the LLM. The longest time recorded in our experiment occurred with a 500-line context size, requiring approximately 70 minutes to complete, which is more than double that for a 100-line context size.
Figure 5: ROC-AUC and Time Taken on fixed context sizes vs. Adaptive Context
Interestingly, applying AC as the context size significantly reduced the time required to complete one round of experiments, making it nearly as fast as using a fixed context size of 80 lines. To investigate this phenomenon, we analyzed the distribution of AC sizes across our 512 data samples, as shown in Figure 6. The majority of AC sizes are relatively small, with a median value of 86.5 lines. Given our observation that elapsed time correlates positively with context size, this smaller median size explains the shorter experimental runtime when using AC. The graph also surges in AC frequency at a size of 300 lines. This anomaly arises because, as described in Section II-B, a default context size of 300 lines was applied when no suitable compound statement could fit within the 500-line limit.
In summary, the AC approach not only provides the best discriminative ability between vulnerable and non-vulnerable lines, but also serves as an economical solution. By reducing GPU memory usage and improving throughput, AC achieves greater computational efficiency while maintaining high detection accuracy, making it a practical choice for real-world applications.
Figure 6: Frequency of Different AC Sizes
@@ III-E RQ4: Line-level Vulnerability Detection
In this section, we compare ANVIL, which does not require any labelled vulnerability data, against three state-of-the-art supervised-learning-based vulnerability detectors: LineVul, LineVD and LLMAO. LineVul [fu_linevul_2022] is a software vulnerability detector that offers both function-level and line-level detection capabilities. It fine-tunes CodeBert [feng_codebert_2020] using the BigVul [fan_cc_2020] vulnerability dataset. The detection process begins at the function level, where LineVul assesses whether a function contains any vulnerabilities. If a function is deemed vulnerable, LineVul then evaluates each line within that function. It does this by aggregating the attention scores from CodeBert’s transformer layers for each token in a line, resulting in a composite attention score for that line. A higher score means the line is more likely to be vulnerable. The intuition is that, for a vulnerability prediction task, lines with higher attention scores are deemed more critical to the model output, suggesting a greater likelihood of being linked to the root cause of the vulnerability.
@@
LineVD [hin_linevd_2022], like LineVul, also supports both function-level and line-level detection. LineVD also uses CodeBert to generate code embeddings for each function and its constituent lines. However, it extends its analysis framework by utilizing Joern [noauthor_joern_nodate] to create a graph representation for each function. This graph, along with the code embeddings, are processed using a graph attention network and a multi-layer perceptron to simultaneously detect vulnerabilities at both the function and line levels. This system is trained on the BigVul dataset, similar to LineVul. A distinctive feature of LineVD is its consistency mechanism: it employs element-wise multiplication within its model to ensure that if a function is classified as non-vulnerable, then all lines within that function are also automatically classified as non-vulnerable.
LLMAO [yang_large_2024], in contrast, is an LLM-based line-level vulnerability detector built on the CodeGen [nijkamp_codegen_2023] series. Unlike LineVul and LineVD, which utilize CodeBERT with 350M parameters, LLMAO employs larger models with 6B and 16B parameters, similar in scale to ANVIL. LLMAO adopts a two-phase approach: first, it uses CodeGen with a left-to-right causal mask to collect the final hidden states for each line of code (LoC) within a context window of 128 lines. These hidden states are then concatenated and passed through a transformer-based bidirectional classification head to compute probabilities for each line, indicating their likelihood of being vulnerable.
While LLMAO was originally trained on the Devign [zhou_devign_2019] dataset, we retrained it using the BigVul dataset to ensure consistency with LineVul and LineVD. This decision was driven by Devign’s limited coverage, as it includes only two open-source projects (FFmpeg and Qemu), which fail to provide sufficient generalizability to real-world scenarios involving diverse and massive codebases. In contrast, BigVul contains data from 4,296 open-source projects, offering a significantly broader and more comprehensive representation. To validate this reasoning, we also conducted the same set of experiments using LLMAO’s original checkpoint from their GitHub repository, trained on Devign. These results, presented in Appendix A, demonstrate that adopting the more diverse BigVul training set improves almost all metrics evaluated in this section.
@@
In this section, all experiments on ANVIL are conducted using the CodeLlama-13B model with the hybrid anomaly scoring function (Equation 1) and Adaptive Context (AC), because these configurations generally yield the best performance as demonstrated in answers to the previous RQs. To ensure fairness and comparability in our evaluation, we will use the same metrics as those employed by our baseline models. We evaluate ANVIL vs. LineVul, LineVD, and LLMAO using two types of metrics: 1) We approach vulnerability detection as a binary classification task and measure the line-level classification performance using the ROC-AUC score. This metric treats each line of code as an independent data sample, assessing the overall classification accuracy without setting a specific threshold. 2) To gauge how effective ANVIL’s anomaly score is as a vulnerability prioritization tool, we evaluate all detectors with a ranking task in which each line within a known vulnerable function is assigned a detector-specific score. We then use Top-N accuracy (a standard metric for measuring vulnerability prioritization ability) to assess whether any of the vulnerable lines are included in the detector’s top-n most confident predictions. Additionally, we report the Mean First Ranking (MFR) to provide an overall measure of the models’ vulnerability prioritization capabilities. For these metrics, each function with known vulnerabilities is treated as a separate data sample.
All evaluations in this section are performed on the PrimeVul dataset, as described in Section III-A. We selected this dataset over Magma and BigVul to mitigate potential overfitting: the supervised-learning-based detectors were trained and tuned on the BigVul dataset, while ANVIL, despite not requiring supervised training, relied on the Magma dataset to select its configurations in previous experiments. These processes could introduce biases specific to their respective datasets. By using PrimeVul as a separate evaluation dataset, we ensure independence and promote a fairer comparison.
TABLE IV: Vulnerability Detection & Localization on PrimeVul Dataset
[…]
ANVIL (7B) N/A 10.3% 26.7% 37.0% 28.9 0.23 55.7%
ANVIL (13B) N/A 9.9% 27.8% 37.4% 26.8 0.22 56.8%
LLMAO (6B) BigVul 7.4% 19.9% 26.6% 50.8 0.40 51.5%
LLMAO (16B) BigVul 9.2% 17.3% 26.2% 49.3 0.39 51.8%
LineVD BigVul 9.2% 15.4% 23.1% 38.0 0.30 49.6%
LineVul BigVul 4.0% 17.0% 30.0% 17.6 0.71 49.6%